OpenStudy (anonymous):

Evaluate without using a calculator ln2+2ln3 - ln 18

OpenStudy (anonymous):

Answer is 0 using a calculator, but i cant this answer without using a calculator

OpenStudy (anonymous):

\[\Large \ln(2)+2\cdot \ln(3)-\ln(18)=\ln(2)+\ln(3^2)-\ln(18)\] \[\Large \ln(2\cdot 3^2)-\ln(18)=\ln(18)-\ln(18)=\] \[\Large \ln\left(\frac{18}{18}\right)=\ln(1)\] and we know that \[\Huge \ln(1)=0\] :)

OpenStudy (anonymous):

there's an equal sign on the second row in the beginning ...

OpenStudy (anonymous):

just want to confirm if it is lnx + lnx u multiply x and x together?

OpenStudy (anonymous):

yes... ln and log have same rules: the difference is: normal... \[\LARGE \log_{10}(x)\] \[\LARGE \ln_{e}(x)\] and when you have: \[\LARGE \ln(x)+\ln(x)=\ln(x\cdot x)=\ln(x^2)\]

OpenStudy (anonymous):

oh ok thanks so much :)